OpenStudy (anonymous):

The open area south of the White House is known as the Ellipse, or President’s Park South. It is 902 feet wide and 1058 ft long. Assume the origin is at the center of the President’s Park South. What is the equation of the ellipse in standard form?

OpenStudy (anonymous):

@jdoe0001 @phi

OpenStudy (anonymous):

@StudyGurl14 can you help me?

OpenStudy (anonymous):

@thomaster @robtobey

OpenStudy (zpupster):

ellipse in standard form x^2/a^2+y^2/b^2=1 a and b are distance to center which would one half 902/2 =a 1058/2=b just plug them in and that is the equation